# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v1

@@ Abstract
With $700$ stars on GitHub and part of the official ONNX repository, the ONNX Optimizer consists of the standard method to apply graph-based optimizations on ONNX models. However, its ability to preserve model accuracy across optimizations, has not been rigorously explored. We propose OODTE, a utility to automatically and thoroughly assess the correctness of the ONNX Optimizer. OODTE follows a simple, yet effective differential testing and evaluation approach that can be easily adopted to other compiler optimizers. In particular, OODTE utilizes a number of ONNX models, then optimizes them and executes both the original and the optimized variants across a user-defined set of inputs, while automatically logging any issues with the optimization process. Finally, for successfully optimized models, OODTE compares the results, and, if any accuracy deviations are observed, it iteratively repeats the process for each pass of the ONNX Optimizer, to localize the root cause of the differences observed. Using OODTE, we sourced well-known $130$ models from the official ONNX Model Hub, used for a wide variety of tasks (classification, object detection, semantic segmentation, text summarization, question and answering, sentiment analysis) from the official ONNX model hub. We observed $9.2$% of all model instances presenting issues leading into the crash of the optimizer, or the generation of an invalid model while using the primary optimizer strategies. In addition, $30$% of the classification models presented accuracy differences across the original and the optimized model variants, while $16.6$% of semantic segmentation and object detection models are also affected, at least to a limited extent. Finally, models related with text analysis and generation were found to be generally robust to the optimization process. Overall, we detected 15 issues, 14 of which were previously unknown, associated with optimizer crashes and accuracy deviations observed in $9$ out of the $47$ available optimizer passes but also the optimizer as a utility in general, which we reported to the developers of the ONNX Optimizer. Our method is simple, yet effective as indicated by the results, and can be easily adapted to other AI model optimization tools and compilers, even beyond the ONNX domain.
@@ I Introduction
Artificial Intelligence has led to a technological revolution, not only within Computer Science but across various aspects of daily life. Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, both researchers and practitioners have developed tools and environments aimed at optimizing DL models [13, 17]. Additionally, widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
@@
Models can also be represented using industry-standard formats, such as ONNX [9]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [5] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question. The ONNX ecosystem also includes additional tools that can further optimize models, potentially at the cost of accuracy—such as the ONNX Quantizer [2]. Developers are typically aware of these trade-offs and apply such tools with consideration of their impact.
@@
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the OONX Optimizer Differential Testing Engine (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, and text generation from the official ONNX Model Hub [8]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
@@
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [6], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. However, our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks. Our methodology builds upon established principles of differential testing [45], which have traditionally been employed for detecting inconsistencies across software variants or transformations.
@@
In total, we observed accuracy deviations in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
@@ II-A Deep Learning Models
Deep Learning Models (DNNs) are a subset of AI models widely used for prediction tasks across various domains. We focus on the following types of models:
@@ II-A1 Image Classification
One of the earliest and most prevalent applications of DNNs is image classification, typically performed using Convolutional Neural Networks (CNNs). These models employ convolution operations, which function as filters to extract specific features from the input image, using weights and biases, of which values are formed during training. Image classification models are often of small-to-medium size in terms of parameters and layers, taking an image (often preprocessed) as input and outputting a list of top-K prediction pairs (prediction, certainty_score). The sum of the $K$ certainty_score instances equals $1$, with the final operation commonly being a Softmax [47] to amplify the highest prediction probability.
II-A2 Object Detection & Semantic Segmentation
Object detection models extend image classification models by not only classifying objects but also detecting where they are placed within an image, typically using bounding boxes. These models also heavily rely on convolution operations. Object detection is widely used in domains like autonomous driving, where multiple objects need to be detected within a scene. A subset of object detection tasks is semantic segmentation, which classifies each pixel in an image into regions corresponding to different object categories. Each pixel is assigned a specific semantic label. For example, a semantic segmentation model for an autonomous vehicle might label parts of a scene as ”road,” ”cars,” ”pedestrians,” etc., providing a more detailed understanding of the environment than simple object detection.
II-A3 Text Generation
Text generation models are a class of models focused on tasks such as text recognition, summarization, sentiment analysis, and more. These models function similarly to image recognition models, receiving an input and generating an output prediction. In text generation, the model typically receives a sequence of natural language input tokens, generates a prediction for the next token, appends it to the input, and discards the leftmost token of the sequence. This process continues iteratively until a stopping condition is met. This approach is referred to as the sliding window technique.
Text generation models are a subset of Generative AI models, responsible for generating content. A relatively new but rapidly evolving family of models in this category are transformers [54], which use an architecture that processes input data to create internal representations, capturing context. These representations are then used to generate new information, typically in the form of text, by predicting one token at a time sequentially (using the sliding window technique described earlier) until the model output is fully generated.
@@ II-B Graph Optimizations
II-C Comparison Methods
Based on the model type under test, OODTE utilizes a set of well-known and established comparison methods. In particular:
  • For classification models, we use Kendall’s Tau Rank Correlation [28], which measures the correlation between ranked lists. This method is suitable for classification tasks, as it compares the order of ranks, yielding a correlation score in the range of $-1$ (completely negative correlation) to $1$ (completely positive correlation). Kendall’s Tau penalizes order discrepancies between the ranks. In classification tasks, each model typically generates a list of top-K predictions per input. We can then extract these top-K predictions from two model variants and apply Kendall’s Tau to compare their outputs.
  • For Object Detection and Semantic Segmentation models, we utilize the following well-known metrics: (1) Average Intersection-over-Union (IoU) – measures the overlap between predicted and ground truth regions, indicating how well a model recognizes class labels. (2) F1 Score – the harmonic mean of precision and recall, balancing false positives and false negatives. (3) Average Precision (AP) – evaluates how many of the model’s positive predictions were actually correct, focusing on precision across different thresholds. (4) Average Recall (AR) – measures how many actual positives the model correctly identified, focusing on recall across thresholds. (5) Mean Average Precision (mAP) – aggregates the model’s performance across multiple thresholds, evaluating both detection and localization accuracy. To account for varying levels of detection confidence, we performed calculations for metrics (2)-(5) using three different IoU thresholds: $0.5$, $0.75$, and $0.9$. These thresholds are necessary to assess model performance at different levels of localization accuracy, where higher thresholds (e.g., $0.9$) require more precise object detection in the image.
  • For text generation models, we utilize BLEU Score [46], which measures the precision of n-gram matches between the generated text and reference text, evaluating the quality of the generated output. The base refers to the output from the original model, and each optimized model’s output is compared to this base to assess similarity.
  • For binary sentiment analysis, we compare the predictions of each sample between the original and optimized models, calculating the percentage difference in predictions to assess the impact of optimization on sentiment classification.
@@ III System Architecture & Methodology
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator Module, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE  with its different modules is shown in Figure 1. OODTE  has the following intuitive workflow: load models, optimize them, execute both the original and the optimized variants across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
@@ III-A Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models in order to get optimized and run. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [8], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
@@ III-B Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where it only applies a specific set of passes to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
@@ III-C Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model variants. It then performs model inference, using ONNX Runtime [3], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [4] package.
In order to be able to utilize large-size datasets and extract results from heavier models while running on low-end devices, the Runner Module supports execution in-parts: by definition in OODTE configuration, a run can be performed in $C$ chuncks, where each chunk consists of $dataset/C$, with $dataset\%C$ also being the final chunk of the run. This applies both for the original and the optimized model, and the results are extracted and propagated to the Metrics Comparison Module, for evaluation and reporting. Following these approach, we were able to conduct our survey on two devices designed for home use without relying on a cluster for results, described in Section V.
@@ III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [68]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [46] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question and answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. The results can then be visualized, but also they are utilized from OODTE in order to localize faults. As aforementioned, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass and repeats the process, reporting how each model behaves when optimized on each pass in isolation. These results are also stored in separate reports. The whole process is iterative for all chunks of input data, for each optimization (if an optimized model is found to be problematic) and for all models under test.

### v2

@@ Abstract
With over 700 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that $9.2$% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, $30$% of classification models and $16.6$% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues—14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
@@ I Introduction
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models [13, 17]. Widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
@@
Models can also be represented using industry-standard formats, such as ONNX [8]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [4] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
@@
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, text comprehension and generation from the official ONNX Model Hub [7]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
@@
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [5], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. Our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
@@
In total, we observed output differences between the original models and their and optimized counterparts in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
@@ II-A Deep Learning Models
Deep neural networks (DNNs) are widely used for prediction tasks across domains. We focus on three common categories:
@@ II-A1 Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function [46].
II-A2 Object Detection and Semantic Segmentation
Object detection extends classification by identifying object locations using bounding boxes. Semantic segmentation further refines this by labeling each pixel with a class (e.g., road, pedestrian), offering dense scene understanding. Both rely heavily on convolutional operations and are critical in areas like autonomous driving.
II-A3 Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers [66], use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
@@ II-B Graph Optimizations
II-C The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify graph structure for compatibility, such as rewrite_input_dtype, which adjusts input data types. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
II-D Model Output Comparison Methods
Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts. In particular:
Classification models:
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation [28], which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection and Semantic Segmentation models:
Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of 0.5, 0.75, and 0.9, with the original model as the reference.
Text generation models:
We convert token outputs to plain text and use the BLEU Score [45] to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
Binary sentiment analysis models:
We measure the percentage of prediction label differences between original and optimized models to evaluate the impact of optimization on classification consistency.
@@ III System Architecture & Methodology
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE with its different modules is shown in Figure 1. It has the following intuitive workflow: load models, optimize them, execute both the original and the optimized versions across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
@@ III-A Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [7], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
@@ III-B Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
@@ III-C Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime [2], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [3] package.
To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts. In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset. Any remaining data not evenly divisible is also included in the final chunk of the run. This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting. By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section V.
@@ III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [67]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [45] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. As mentioned earlier, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.